# A Comparison-Relationship-Surrogate Evolutionary Algorithm for Multi-Objective Optimization

### v1

@@ 3.1.2 Crowding Distance
With the ability to calculate the non-domination rank of individuals using the surrogate model, the other problem is the calculation of crowding distance. In the original algorithm, each individual is first assigned a crowding distance of zero. Then, for each of the objective functions, the following operations are performed. Individuals are sorted by the value of that objective function and the solutions with the highest and lowest values are assigned an infinite crowding distance. The crowding distance of the remaining individuals are incremented by the distance between their neighbors in the sorted list (normalized by the full range of the objective functions). This final step involving the distance between solutions cannot be achieved with a model of the comparison relationships.
Although the comparison surrogate model cannot evaluate the crowding distance directly, it can still be used to sort the individuals and find the most extreme examples. We suggest using the values of $\tau$ from Eq. 3 to find these extreme examples and set their crowding distance to infinity. The remaining individuals with a crowding distance of zero are shuffled before being returned.
We have shown how the model of comparison relationships can be used to calculate non-domination rank and a (greatly simplified) version of crowding distance for a given population. From there, the individuals can be sorted by fitness and selected to continue to the next generation. These fitness values can also be used during the creation of children with, for example, tournament selection.
@@ 3.2 Modeling the Comparison Relationship
One way to approach the task of modeling the comparison relationships between objectives is by considering it as multiple binary classification problems. The training data here is the set of pairs of individuals in the evaluated population with the labels being the $m$ comparisons of objective functions ($f_{i}(x_{1})<f_{i}(x_{2})$) for each example. We are interested in building a model from the data to predict the comparisons for the unseen pairs of individuals. Due to their expressive power and success in solving problems with high dimensional inputs, we used a deep neural network to approximate the comparison relationship.
@@
Figure 2: A schematic of the comparison surrogate model. Decision variables of the two individuals being compared are fed into feature extraction networks made of densely connected layers with shared weights. Additional features in the form of the comparisons $x_{1,i}<x_{2,i}$ are generated. All features are combined into a symmetric vector such that a permutation of the two input individuals will cause the feature vector to reverse itself ($f_{i}\to f_{n-i+1}$ for a vector $f$ with $n$ elements). These features are fed into comparison networks (one for each objective) which use special symmetric dense layers to ensure the correct behavior under permutation of the inputs.
Considering a network that outputs predictions with the one-hot encoding scheme, the requirement in Eq. 4 is equivalent to requiring that the output vector of the network is reversed on permutation of the two inputs. To achieve this, the decision variables for the two individuals being compared are each first fed into a feature extraction network that consists of conventional dense layers with ReLU activation and has shared weights and biases between the two inputs. The output features of the two individuals are concatenated into a single feature vector with the elements from the second network being reversed so that permuting the individuals is equivalent to reversing the full feature vector. That is,
@@
Figure 3: An example of the comparison model trained on data from the WFG8 ($n=16$) problem. (a) loss function on the train and test set during training; (b) the accuracy during training; (c) and (d) predicted comparisons for both objectives on never-before-seen samples where individuals are indexed in sorted order for the objective being compared (ie predictions from a perfect classifier would show up as a solid yellow upper triangle); (e) the receiver operating characteristic curve for each model output as well as domination predictions.
@@
An example of using the comparison-relationship-surrogate model is shown in Fig. 3. A training set of 300 individuals was generated for the problem WFG8 ($n=16$[44] and the model was trained on it for 128 epochs. The resulting learning curves are shown in subfigures (a) and (b) using an 80%-20% train-test split of the labeled comparison data. Predicted comparisons were then evaluated for every pair in a second set of 128 randomly generated individuals. The results are plotted in subfigures (c) and (d) where the individuals are indexed according to their position in the array sorted by the objective being evaluated. This representation helps visualize when and where errors occur in the comparisons. For example, many of the out-of-place elements are close to the diagonal meaning that errors tend to be between individuals that are similar in value for that objective. In subfigure (e) we show the receiver operating characteristic curve and list the area under curve metric for the comparison models as well as for the resulting domination classifier. As expected, the domination predictions are degraded as a correct prediction requires that both underlying comparisons are predicted correctly.
@@ 4 Empirical Studies of Performance
Three different classes of algorithms were chosen to compare against: conventional genetic algorithms, regression-based surrogate-assisted algorithms, and classification-based surrogate-assisted algorithms. From the first class, NSGA-II [38] was chosen for test problems with two objectives. We feel this is the best choice as the new comparison-surrogate-assisted algorithm is directly based on NSGA-II. For the problems with more than two objectives, NSGA-III [46] was chosen as the conventional optimizer due to its improved performance on many-objective problems. From the category of regression-based surrogate-assisted optimizers, we chose ParEGO [47], K-RVEA [11], and MOEA/D-EGO [48] for a combination of their performance on test problems and popularity in the literature. Because the new algorithm being studied uses a classification-based surrogate model, we feel it is important to include similar algorithms in the numerical experiments. For these purposes, we chose an example of an algorithm that classifies the regions of decision space, CPS-MOEA [16], as well as one that uses a pairwise model of the domination relationships, REMO [25].
All algorithms except for the new comparison-relationship-surrogate-based optimizer were evaluated using their implementation in PlatEMO v4.1 [49] with default settings for all parameters. For algorithms that require a population size, it was set to 50 in all cases, except for NSGA-II where it was set to 20 following the recommendation of Ref. [47] and NSGA-III where it was set based on the number of objectives using the table in Ref. [23]. For surrogate-assisted algorithms that use an initial population size, it was set to $N_{i}=11n-1$ as recommended for the Latin hypercube sampling technique [37] and was capped at $N_{i}=100$ for problems with more than ten decision variables to avoid immediately exhausting the budget of function evaluations.
@@
Problem m NSGA-III ParEGO K-RVEA MOEA/D-EGO CPS-MOEA REMO CRSEA DTLZ1 3 9.09e+1 (1.57e+1) - 6.73e+1 (5.57e+0) + 9.01e+1 (2.27e+1) - 8.43e+1 (1.62e+1) $\approx$ 8.00e+1 (1.57e+1) $\approx$ 5.40e+1 (1.87e+1) + 8.11e+1 (1.65e+1) 4 6.04e+1 (1.62e+1) $\approx$ 5.63e+1 (7.24e+0) + 6.23e+1 (1.41e+1) $\approx$ 6.21e+1 (1.23e+1) $\approx$ 5.94e+1 (1.35e+1) $\approx$ 3.78e+1 (1.20e+1) + 6.26e+1 (1.23e+1) 6 3.11e+1 (1.07e+1) + 3.71e+1 (5.59e+0) + 2.94e+1 (7.54e+0) + 3.32e+1 (7.47e+0) + 3.09e+1 (8.56e+0) + 1.53e+1 (4.56e+0) + 4.21e+1 (7.12e+0) 8 1.08e+1 (5.16e+0) + 1.55e+1 (4.15e+0) $\approx$ 9.37e+0 (3.43e+0) + 1.28e+1 (3.27e+0) + 1.01e+1 (3.56e+0) + 4.55e+0 (2.28e+0) + 1.53e+1 (5.47e+0) 10 5.22e-1 (2.29e-1) + 4.58e-1 (8.84e-2) + 3.68e-1 (1.07e-1) + 4.61e-1 (1.62e-1) + 3.99e-1 (8.58e-2) + 2.95e-1 (6.39e-2) + 6.91e-1 (2.99e-1) DTLZ2 3 2.93e-1 (2.69e-2) - 1.60e-1 (3.75e-2) - 1.73e-1 (4.10e-2) - 3.54e-1 (2.27e-2) - 2.86e-1 (2.10e-2) - 2.09e-1 (2.56e-2) - 8.36e-2 (1.03e-2) 4 3.52e-1 (2.95e-2) - 4.28e-1 (3.04e-2) - 2.41e-1 (3.13e-2) - 3.95e-1 (2.07e-2) - 3.95e-1 (3.16e-2) - 2.71e-1 (2.33e-2) - 2.17e-1 (2.44e-2) 6 4.83e-1 (3.55e-2) $\approx$ 5.33e-1 (2.92e-2) - 3.37e-1 (2.38e-2) + 4.81e-1 (1.72e-2) $\approx$ 5.36e-1 (2.93e-2) - 4.04e-1 (4.00e-2) + 4.85e-1 (2.62e-2) 8 5.83e-1 (2.78e-2) - 6.59e-1 (2.77e-2) - 4.67e-1 (3.70e-2) + 5.18e-1 (1.50e-2) + 5.98e-1 (2.90e-2) - 5.13e-1 (4.67e-2) + 5.31e-1 (1.87e-2) 10 6.71e-1 (2.99e-2) - 7.00e-1 (2.87e-2) - 5.41e-1 (2.32e-2) $\approx$ 5.69e-1 (3.21e-2) - 6.55e-1 (3.95e-2) - 6.12e-1 (4.50e-2) - 5.40e-1 (1.71e-2) DTLZ3 3 2.65e+2 (6.68e+1) - 1.75e+2 (8.48e+0) + 2.24e+2 (5.49e+1) $\approx$ 2.03e+2 (3.32e+1) + 2.13e+2 (3.41e+1) $\approx$ 1.63e+2 (4.38e+1) + 2.17e+2 (2.71e+1) 4 1.82e+2 (5.01e+1) $\approx$ 1.50e+2 (9.92e+0) + 1.73e+2 (2.97e+1) $\approx$ 1.66e+2 (2.30e+1) $\approx$ 1.58e+2 (3.09e+1) + 1.24e+2 (2.88e+1) + 1.72e+2 (2.00e+1) 6 9.47e+1 (3.70e+1) + 9.64e+1 (1.23e+1) + 8.25e+1 (2.29e+1) + 9.53e+1 (1.59e+1) + 9.32e+1 (2.06e+1) + 5.60e+1 (1.81e+1) + 1.13e+2 (1.04e+1) 8 3.27e+1 (1.27e+1) + 4.61e+1 (1.00e+1) $\approx$ 2.80e+1 (1.07e+1) + 3.62e+1 (1.23e+1) + 3.41e+1 (1.33e+1) + 1.87e+1 (8.97e+0) + 4.52e+1 (1.13e+1) 10 2.02e+0 (1.75e+0) + 1.58e+0 (4.97e-1) + 1.22e+0 (3.10e-1) + 1.43e+0 (4.37e-1) + 2.20e+0 (1.36e+0) $\approx$ 9.63e-1 (2.68e-1) + 1.90e+0 (6.49e-1) DTLZ4 3 7.34e-1 (1.92e-1) $\approx$ 5.31e-1 (1.31e-1) + 4.15e-1 (1.19e-1) + 6.01e-1 (6.79e-2) + 5.62e-1 (6.06e-2) + 2.62e-1 (1.27e-1) + 7.26e-1 (1.71e-1) 4 7.36e-1 (1.60e-1) $\approx$ 6.93e-1 (1.02e-1) $\approx$ 5.12e-1 (9.17e-2) + 6.55e-1 (5.16e-2) $\approx$ 6.16e-1 (3.91e-2) + 3.22e-1 (7.95e-2) + 7.06e-1 (1.96e-1) 6 7.74e-1 (9.65e-2) - 5.00e-1 (7.57e-2) + 5.42e-1 (4.58e-2) + 6.54e-1 (3.23e-2) - 6.34e-1 (2.36e-2) - 5.10e-1 (4.82e-2) + 6.05e-1 (5.73e-2) 8 7.54e-1 (6.06e-2) - 5.50e-1 (3.76e-2) + 5.62e-1 (3.80e-2) + 6.29e-1 (1.83e-2) $\approx$ 6.31e-1 (1.36e-2) $\approx$ 5.92e-1 (4.34e-2) + 6.31e-1 (4.40e-2) 10 7.15e-1 (6.09e-2) - 5.97e-1 (2.37e-2) - 6.09e-1 (2.62e-2) - 6.13e-1 (1.16e-2) - 6.19e-1 (8.78e-3) - 6.43e-1 (2.86e-2) - 5.84e-1 (2.03e-2) DTLZ5 3 2.03e-1 (3.90e-2) - 5.63e-2 (4.49e-2) - 8.68e-2 (2.19e-2) - 2.52e-1 (3.16e-2) - 2.14e-1 (3.31e-2) - 1.00e-1 (2.86e-2) - 4.09e-2 (1.58e-2) 4 1.89e-1 (2.61e-2) - 1.06e-1 (6.76e-2) + 6.11e-2 (2.28e-2) + 2.25e-1 (2.87e-2) - 1.95e-1 (3.11e-2) - 1.19e-1 (2.52e-2) $\approx$ 1.10e-1 (2.49e-2) 6 1.35e-1 (3.17e-2) - 1.03e-1 (3.61e-2) + 5.23e-2 (1.20e-2) + 1.41e-1 (1.59e-2) - 1.46e-1 (2.54e-2) - 6.45e-2 (1.99e-2) + 1.20e-1 (2.37e-2) 8 8.73e-2 (1.44e-2) - 7.40e-2 (9.77e-3) - 2.37e-2 (5.86e-3) + 6.85e-2 (7.06e-3) $\approx$ 7.43e-2 (1.12e-2) - 3.45e-2 (9.26e-3) + 6.89e-2 (1.02e-2) 10 2.95e-2 (5.14e-3) - 1.94e-2 (1.73e-3) $\approx$ 1.25e-2 (2.08e-3) + 1.94e-2 (2.52e-3) $\approx$ 1.90e-2 (2.53e-3) $\approx$ 1.22e-2 (2.51e-3) + 2.01e-2 (3.68e-3) DTLZ6 3 6.08e+0 (3.33e-1) - 9.22e-1 (1.87e-1) + 2.88e+0 (4.35e-1) $\approx$ 2.21e+0 (8.20e-1) + 3.48e+0 (5.78e-1) - 4.30e+0 (6.05e-1) - 2.77e+0 (9.65e-1) 4 5.31e+0 (3.56e-1) + 9.95e-1 (1.77e-1) + 1.97e+0 (3.26e-1) + 1.71e+0 (5.97e-1) + 3.21e+0 (6.65e-1) + 4.10e+0 (7.11e-1) + 5.44e+0 (3.04e-1) 6 3.76e+0 (2.91e-1) + 6.31e-1 (1.26e-1) + 1.50e+0 (3.49e-1) + 1.20e+0 (4.96e-1) + 2.20e+0 (5.10e-1) + 2.27e+0 (6.98e-1) + 4.07e+0 (1.24e-1) 8 2.14e+0 (2.80e-1) + 4.35e-1 (1.12e-1) + 5.13e-1 (2.30e-1) + 4.92e-1 (2.18e-1) + 1.04e+0 (3.77e-1) + 8.31e-1 (4.07e-1) + 2.37e+0 (9.13e-2) 10 5.31e-1 (2.14e-1) + 1.44e-1 (5.45e-2) + 6.25e-2 (2.98e-2) + 2.21e-1 (8.24e-2) + 6.70e-2 (3.34e-2) + 8.50e-2 (9.64e-2) + 7.01e-1 (3.69e-2) DTLZ7 3 4.87e+0 (8.89e-1) - 1.51e-1 (1.48e-2) + 1.29e-1 (1.90e-2) + 3.04e-1 (1.04e-1) + 3.73e+0 (1.30e+0) - 8.11e-1 (5.20e-1) - 5.11e-1 (2.17e-1) 4 5.66e+0 (1.41e+0) - 3.04e-1 (3.35e-2) + 3.15e-1 (3.18e-2) + 5.32e-1 (5.53e-2) + 4.49e+0 (2.12e+0) - 1.27e+0 (4.78e-1) - 7.12e-1 (2.63e-1) 6 8.57e+0 (2.66e+0) - 7.48e-1 (3.70e-2) + 6.17e-1 (2.84e-2) + 9.28e-1 (7.24e-2) + 4.23e+0 (2.51e+0) - 2.86e+0 (4.69e-1) - 1.10e+0 (1.13e-1) 8 8.80e+0 (3.16e+0) - 1.18e+0 (4.96e-2) + 9.63e-1 (6.06e-2) + 1.12e+0 (4.02e-2) + 2.79e+0 (1.94e+0) - 3.53e+0 (5.02e-1) - 1.77e+0 (2.09e-1) 10 3.03e+0 (1.51e+0) - 1.41e+0 (5.75e-2) + 1.05e+0 (2.92e-2) + 1.26e+0 (3.21e-2) + 1.56e+0 (1.09e-1) + 2.21e+0 (3.91e-1) - 1.89e+0 (2.11e-1) +/-$\approx$ 10/20/5 23/8/4 25/5/5 19/8/8 13/16/6 23/11/1

### v2

@@ 3.2.1 Crowding Distance
With the ability to clean the model’s predictions from violations of the transitive property, we can now use it to predict the non-domination rank of solutions. In NSGA-II, “crowding distance” is then used to break ties of non-domination rank for the purpose of selection. This acts as a form of diversity preservation.
To calculate crowding distance, each solution is assigned a value of zero to start with. For each objective, the solutions are sorted by the value of their objective function. The extreme solutions (highest and lowest value) are assigned a crowding distance of infinity. Every other solution has its crowding distance incremented by the distance between its two nearest neighbors for this objective normalized by full range spanned by all solutions (distance between the biggest and smallest values).
The comparison-relationship model, of course, does not have information on the distance between neighbors. Instead we suggest using the values of $\tau$ from Eq. 2 to predict the solutions with extreme values for the objective functions (smallest and largest). These solutions are preferred when breaking ties of non-domination rank. This should achieve a similar effect to setting their crowding distance to infinity in NSGA-II, but we do expect worse diversity preservation since the surrogate model cannot predict the distance between solutions.
With these two adjustments (cleaning the predictions to satisfy the transitive property, and breaking ties with predictions of extreme solutions instead of crowding distance) the comparison-relationship model may be used in an NSGA-II-like algorithm to perform selection.
@@ 3.3 Modeling the Comparison Relationship
One way to approach the task of modeling the comparison relationships between objectives is by considering it as multiple binary classification problems. The training data here is the set of pairs of solutions in the evaluated population with the labels being the $m$ comparisons of objective functions ($f_{i}(x_{1})<f_{i}(x_{2})$) for each example. We are interested in building a model from the data to predict the comparisons for the unseen pairs of solutions. Due to their expressive power and success in solving problems with high dimensional inputs, we used a deep neural network to approximate the comparison relationship.
@@
Figure 2: A schematic of the comparison surrogate model. Decision variables of the two solutions being compared are fed into feature extraction networks made of densely connected layers with shared weights. Additional features in the form of the comparisons $x_{1,i}<x_{2,i}$ are generated. All features are combined into a symmetric vector such that a permutation of the two input solutions will cause the feature vector to reverse itself ($f_{i}\to f_{n-i+1}$ for a vector $f$ with $n$ elements). These features are fed into comparison networks (one for each objective) which use special symmetric dense layers to ensure the correct behavior under permutation of the inputs.
Considering a network that outputs predictions with the one-hot encoding scheme, the requirement in Eq. 3 is equivalent to requiring that the output vector of the network is reversed on permutation of the two inputs. To achieve this, the decision variables for the two solutions being compared are each first fed into a feature extraction network that consists of conventional dense layers with ReLU activation and has shared weights and biases between the two inputs. The output features of the two solutions are concatenated into a single feature vector with the elements from the second network being reversed so that permuting the solutions is equivalent to reversing the full feature vector. That is,
@@
Figure 3: An example of the comparison model trained on data from the WFG8 ($n=16$) problem. (a) loss function on the train and test set during training; (b) the accuracy during training; (c) and (d) predicted comparisons for both objectives on never-before-seen samples where solutions are indexed in sorted order for the objective being compared (i.e. predictions from a perfect classifier would show up as a solid yellow upper triangle); (e) the receiver operating characteristic curve for each model output as well as domination predictions.
@@
An example of using the comparison-relationship-surrogate model is shown in Fig. 3. A training set of 300 solutions was generated for the problem WFG8 ($n=16$[48] and the model was trained on it for 128 epochs. The resulting learning curves are shown in subfigures (a) and (b) using an 80%-20% train-test split of the labeled comparison data. Predicted comparisons were then evaluated for every pair in a second set of 128 randomly generated solutions. The results are plotted in subfigures (c) and (d) where the solutions are indexed according to their position in the array sorted by the objective being evaluated. This representation helps visualize when and where errors occur in the comparisons. For example, many of the out-of-place elements are close to the diagonal meaning that errors tend to be between solutions that are similar in value for that objective. In subfigure (e) we show the receiver operating characteristic curve and list the area under curve metric for the comparison models as well as for the resulting domination classifier. As expected, the domination predictions are degraded as a correct prediction requires that both underlying comparisons are predicted correctly.
@@ 4 Empirical Studies of Performance
Three different classes of algorithms were chosen to compare against: conventional genetic algorithms, regression-based surrogate-assisted algorithms, and classification-based surrogate-assisted algorithms. From the first class, NSGA-II [42] was chosen for test problems with two objectives. This algorithm was selected for its popularity in applications as well as the fact that the new comparison-surrogate-assisted algorithm is strongly influenced by it. For the problems with more than two objectives, NSGA-III [50] was chosen as the conventional optimizer due to its improved performance on many-objective problems. From the category of regression-based surrogate-assisted optimizers, we chose ParEGO [14], K-RVEA [11], and MOEA/D-EGO [17] for a combination of their performance on test problems and popularity in the literature. Because the new algorithm being studied uses a classification-based surrogate model, we feel it is important to include similar algorithms in the numerical experiments. For these purposes, we chose an example of an algorithm that classifies the regions of decision space, CPS-MOEA [20], as well as two that use a pairwise model of the domination relationships, REMO [29] and PC-SAEA [32].
All algorithms except for the new comparison-relationship-surrogate-based optimizer were evaluated using their implementation in PlatEMO [51] version 4.11 with default settings for all parameters. For algorithms that require a population size, it was set to 50 in all cases, except for NSGA-II where it was set to 20 following the recommendation of Ref. [14] and NSGA-III where it was set based on the number of objectives using the table in Ref. [27]. For surrogate-assisted algorithms that use an initial population size, it was set to $N_{i}=11n-1$ as recommended for the Latin hypercube sampling technique [41] and was capped at $N_{i}=100$ for problems with more than ten decision variables to avoid immediately exhausting the budget of function evaluations.
@@
Problem m NSGA-III ParEGO K-RVEA MOEA/D-EGO CPS-MOEA REMO PC-SAEA CRSEA DTLZ1 3 9.09e+1 (1.57e+1) - 6.73e+1 (5.57e+0) + 9.01e+1 (2.27e+1) - 8.43e+1 (1.62e+1) $\approx$ 8.00e+1 (1.57e+1) $\approx$ 5.40e+1 (1.87e+1) + 8.61e+1 (1.66e+1) $\approx$ 8.11e+1 (1.65e+1) 4 6.04e+1 (1.62e+1) $\approx$ 5.63e+1 (7.24e+0) + 6.23e+1 (1.41e+1) $\approx$ 6.21e+1 (1.23e+1) $\approx$ 5.94e+1 (1.35e+1) $\approx$ 3.78e+1 (1.20e+1) + 7.03e+1 (1.60e+1) - 6.26e+1 (1.23e+1) 6 3.11e+1 (1.07e+1) + 3.71e+1 (5.59e+0) + 2.94e+1 (7.54e+0) + 3.32e+1 (7.47e+0) + 3.09e+1 (8.56e+0) + 1.53e+1 (4.56e+0) + 2.76e+1 (1.12e+1) + 4.21e+1 (7.12e+0) 8 1.08e+1 (5.16e+0) + 1.55e+1 (4.15e+0) $\approx$ 9.37e+0 (3.43e+0) + 1.28e+1 (3.27e+0) + 1.01e+1 (3.56e+0) + 4.55e+0 (2.28e+0) + 9.16e+0 (3.17e+0) + 1.53e+1 (5.47e+0) 10 5.22e-1 (2.29e-1) + 4.58e-1 (8.84e-2) + 3.68e-1 (1.07e-1) + 4.61e-1 (1.62e-1) + 3.99e-1 (8.58e-2) + 2.95e-1 (6.39e-2) + 3.91e-1 (8.21e-2) + 6.91e-1 (2.99e-1) DTLZ2 3 2.93e-1 (2.69e-2) - 1.60e-1 (3.75e-2) - 1.73e-1 (4.10e-2) - 3.54e-1 (2.27e-2) - 2.86e-1 (2.10e-2) - 2.09e-1 (2.56e-2) - 2.80e-1 (2.39e-2) - 8.36e-2 (1.03e-2) 4 3.52e-1 (2.95e-2) - 4.28e-1 (3.04e-2) - 2.41e-1 (3.13e-2) - 3.95e-1 (2.07e-2) - 3.95e-1 (3.16e-2) - 2.71e-1 (2.33e-2) - 3.51e-1 (2.01e-2) - 2.17e-1 (2.44e-2) 6 4.83e-1 (3.55e-2) $\approx$ 5.33e-1 (2.92e-2) - 3.37e-1 (2.38e-2) + 4.81e-1 (1.72e-2) $\approx$ 5.36e-1 (2.93e-2) - 4.04e-1 (4.00e-2) + 4.69e-1 (2.70e-2) + 4.85e-1 (2.62e-2) 8 5.83e-1 (2.78e-2) - 6.59e-1 (2.77e-2) - 4.67e-1 (3.70e-2) + 5.18e-1 (1.50e-2) + 5.98e-1 (2.90e-2) - 5.13e-1 (4.67e-2) + 5.71e-1 (2.07e-2) - 5.31e-1 (1.87e-2) 10 6.71e-1 (2.99e-2) - 7.00e-1 (2.87e-2) - 5.41e-1 (2.32e-2) $\approx$ 5.69e-1 (3.21e-2) - 6.55e-1 (3.95e-2) - 6.12e-1 (4.50e-2) - 6.72e-1 (2.91e-2) - 5.40e-1 (1.71e-2) DTLZ3 3 2.65e+2 (6.68e+1) - 1.75e+2 (8.48e+0) + 2.24e+2 (5.49e+1) $\approx$ 2.03e+2 (3.32e+1) + 2.13e+2 (3.41e+1) $\approx$ 1.63e+2 (4.38e+1) + 2.38e+2 (4.10e+1) - 2.17e+2 (2.71e+1) 4 1.82e+2 (5.01e+1) $\approx$ 1.50e+2 (9.92e+0) + 1.73e+2 (2.97e+1) $\approx$ 1.66e+2 (2.30e+1) $\approx$ 1.58e+2 (3.09e+1) + 1.24e+2 (2.88e+1) + 2.01e+2 (3.98e+1) - 1.72e+2 (2.00e+1) 6 9.47e+1 (3.70e+1) + 9.64e+1 (1.23e+1) + 8.25e+1 (2.29e+1) + 9.53e+1 (1.59e+1) + 9.32e+1 (2.06e+1) + 5.60e+1 (1.81e+1) + 1.01e+2 (2.66e+1) + 1.13e+2 (1.04e+1) 8 3.27e+1 (1.27e+1) + 4.61e+1 (1.00e+1) $\approx$ 2.80e+1 (1.07e+1) + 3.62e+1 (1.23e+1) + 3.41e+1 (1.33e+1) + 1.87e+1 (8.97e+0) + 2.55e+1 (8.64e+0) + 4.52e+1 (1.13e+1) 10 2.02e+0 (1.75e+0) + 1.58e+0 (4.97e-1) + 1.22e+0 (3.10e-1) + 1.43e+0 (4.37e-1) + 2.20e+0 (1.36e+0) $\approx$ 9.63e-1 (2.68e-1) + 1.29e+0 (2.72e-1) + 1.90e+0 (6.49e-1) DTLZ4 3 7.34e-1 (1.92e-1) $\approx$ 5.31e-1 (1.31e-1) + 4.15e-1 (1.19e-1) + 6.01e-1 (6.79e-2) + 5.62e-1 (6.06e-2) + 2.62e-1 (1.27e-1) + 5.26e-1 (8.64e-2) + 7.26e-1 (1.71e-1) 4 7.36e-1 (1.60e-1) $\approx$ 6.93e-1 (1.02e-1) $\approx$ 5.12e-1 (9.17e-2) + 6.55e-1 (5.16e-2) $\approx$ 6.16e-1 (3.91e-2) + 3.22e-1 (7.95e-2) + 6.01e-1 (5.30e-2) + 7.06e-1 (1.96e-1) 6 7.74e-1 (9.65e-2) - 5.00e-1 (7.57e-2) + 5.42e-1 (4.58e-2) + 6.54e-1 (3.23e-2) - 6.34e-1 (2.36e-2) - 5.10e-1 (4.82e-2) + 7.06e-1 (3.64e-2) - 6.05e-1 (5.73e-2) 8 7.54e-1 (6.06e-2) - 5.50e-1 (3.76e-2) + 5.62e-1 (3.80e-2) + 6.29e-1 (1.83e-2) $\approx$ 6.31e-1 (1.36e-2) $\approx$ 5.92e-1 (4.34e-2) + 6.62e-1 (1.98e-2) - 6.31e-1 (4.40e-2) 10 7.15e-1 (6.09e-2) - 5.97e-1 (2.37e-2) - 6.09e-1 (2.62e-2) - 6.13e-1 (1.16e-2) - 6.19e-1 (8.78e-3) - 6.43e-1 (2.86e-2) - 6.35e-1 (2.13e-2) - 5.84e-1 (2.03e-2) DTLZ5 3 2.03e-1 (3.90e-2) - 5.63e-2 (4.49e-2) - 8.68e-2 (2.19e-2) - 2.52e-1 (3.16e-2) - 2.14e-1 (3.31e-2) - 1.00e-1 (2.86e-2) - 2.06e-1 (2.98e-2) - 4.09e-2 (1.58e-2) 4 1.89e-1 (2.61e-2) - 1.06e-1 (6.76e-2) + 6.11e-2 (2.28e-2) + 2.25e-1 (2.87e-2) - 1.95e-1 (3.11e-2) - 1.19e-1 (2.52e-2) $\approx$ 1.89e-1 (2.26e-2) - 1.10e-1 (2.49e-2) 6 1.35e-1 (3.17e-2) - 1.03e-1 (3.61e-2) + 5.23e-2 (1.20e-2) + 1.41e-1 (1.59e-2) - 1.46e-1 (2.54e-2) - 6.45e-2 (1.99e-2) + 1.18e-1 (1.95e-2) $\approx$ 1.20e-1 (2.37e-2) 8 8.73e-2 (1.44e-2) - 7.40e-2 (9.77e-3) - 2.37e-2 (5.86e-3) + 6.85e-2 (7.06e-3) $\approx$ 7.43e-2 (1.12e-2) - 3.45e-2 (9.26e-3) + 5.50e-2 (5.78e-3) + 6.89e-2 (1.02e-2) 10 2.95e-2 (5.14e-3) - 1.94e-2 (1.73e-3) $\approx$ 1.25e-2 (2.08e-3) + 1.94e-2 (2.52e-3) $\approx$ 1.90e-2 (2.53e-3) $\approx$ 1.22e-2 (2.51e-3) + 1.39e-2 (1.56e-3) + 2.01e-2 (3.68e-3) DTLZ6 3 6.08e+0 (3.33e-1) - 9.22e-1 (1.87e-1) + 2.88e+0 (4.35e-1) $\approx$ 2.21e+0 (8.20e-1) + 3.48e+0 (5.78e-1) - 4.30e+0 (6.05e-1) - 6.06e+0 (3.90e-1) - 2.77e+0 (9.65e-1) 4 5.31e+0 (3.56e-1) + 9.95e-1 (1.77e-1) + 1.97e+0 (3.26e-1) + 1.71e+0 (5.97e-1) + 3.21e+0 (6.65e-1) + 4.10e+0 (7.11e-1) + 5.24e+0 (2.20e-1) + 5.44e+0 (3.04e-1) 6 3.76e+0 (2.91e-1) + 6.31e-1 (1.26e-1) + 1.50e+0 (3.49e-1) + 1.20e+0 (4.96e-1) + 2.20e+0 (5.10e-1) + 2.27e+0 (6.98e-1) + 3.64e+0 (2.37e-1) + 4.07e+0 (1.24e-1) 8 2.14e+0 (2.80e-1) + 4.35e-1 (1.12e-1) + 5.13e-1 (2.30e-1) + 4.92e-1 (2.18e-1) + 1.04e+0 (3.77e-1) + 8.31e-1 (4.07e-1) + 1.89e+0 (1.76e-1) + 2.37e+0 (9.13e-2) 10 5.31e-1 (2.14e-1) + 1.44e-1 (5.45e-2) + 6.25e-2 (2.98e-2) + 2.21e-1 (8.24e-2) + 6.70e-2 (3.34e-2) + 8.50e-2 (9.64e-2) + 4.82e-1 (1.98e-1) + 7.01e-1 (3.69e-2) DTLZ7 3 4.87e+0 (8.89e-1) - 1.51e-1 (1.48e-2) + 1.29e-1 (1.90e-2) + 3.04e-1 (1.04e-1) + 3.73e+0 (1.30e+0) - 8.11e-1 (5.20e-1) - 4.78e+0 (1.13e+0) - 5.11e-1 (2.17e-1) 4 5.66e+0 (1.41e+0) - 3.04e-1 (3.35e-2) + 3.15e-1 (3.18e-2) + 5.32e-1 (5.53e-2) + 4.49e+0 (2.12e+0) - 1.27e+0 (4.78e-1) - 6.25e+0 (1.65e+0) - 7.12e-1 (2.63e-1) 6 8.57e+0 (2.66e+0) - 7.48e-1 (3.70e-2) + 6.17e-1 (2.84e-2) + 9.28e-1 (7.24e-2) + 4.23e+0 (2.51e+0) - 2.86e+0 (4.69e-1) - 8.03e+0 (2.52e+0) - 1.10e+0 (1.13e-1) 8 8.80e+0 (3.16e+0) - 1.18e+0 (4.96e-2) + 9.63e-1 (6.06e-2) + 1.12e+0 (4.02e-2) + 2.79e+0 (1.94e+0) - 3.53e+0 (5.02e-1) - 6.35e+0 (2.83e+0) - 1.77e+0 (2.09e-1) 10 3.03e+0 (1.51e+0) - 1.41e+0 (5.75e-2) + 1.05e+0 (2.92e-2) + 1.26e+0 (3.21e-2) + 1.56e+0 (1.09e-1) + 2.21e+0 (3.91e-1) - 1.79e+0 (2.79e-1) + 1.89e+0 (2.11e-1) +/-$\approx$ 10/20/5 23/8/4 25/5/5 19/8/8 13/16/6 23/11/1 16/17/2